# Thawing Dark Energy and Massive Neutrinos in Light of DESI

### v2

@@ Abstract
Recent analyses have shown that a dynamic dark energy modeled by the CPL parameterization of the dark energy equation of state (EoS) can ease constraints on the total neutrino mass compared to the standard $\Lambda$CDM model. This helps reconcile cosmological and particle physics measurements of $\sum m_{\nu}$. In this study, we investigate the robustness of this effect by assessing the extent to which the CPL assumption influences the results. We examine how alternative EoS parameterizations - such as Barboza-Alcaniz (BA), Jassal-Bagla-Padmanabhan (JBP), and a physically motivated thawing parameterization that reproduces the behavior of various scalar field models - affect estimates of $\sum m_{\nu}$. Although both the BA and JBP parameterizations relax the constraints similarly to the CPL model, the JBP parameterization still excludes the inverted neutrino mass hierarchy at $\sim 2.1\;\sigma$ with $\sum m_{\nu}<0.096$ eV. The thawing parameterization excludes the inverted hierarchy at $\sim 3.3\sigma$ and yields tighter constraints, comparable to those of the $\Lambda$CDM model, with $\sum m_{\nu}<0.071$ eV. Our analysis also reveals that the preference for unphysical negative neutrino masses is significantly reduced in all models considered. Finally, we show that the thawing model can be mapped into the BA and JBP $w_{0}$$w_{a}$ parameter space, with the apparent preference for the phantom regime actually supporting quintessence (non-phantom) models.
@@ I Introduction
In the constext of General Relativity, two major frameworks have emerged as alternative to model the Universe’s accelerated expansion (see e.g. [5, 6, 7, 8, 9, 10, 11]): the scalar field approach, exemplified by quintessence models, and phenomenological parameterizations of the dark energy equation of state (EoS), such as the widely used Chevallier–Polarski–Linder (CPL), Barboza–Alcaniz (BA), and Jassal–Bagla–Padmanabhan (JBP) forms [12, 13, 14, 15, 16]. In both frameworks, the dark energy EoS is a dynamical quantity, $w(z)$.
@@
Beyond the DDE discussion, the late-time universe is plagued by further uncertainties related to the background and perturbations evolution, and the neutrino physics is of particular interest in this picture. In the standard scenario, outlined by the Standard Model of fundamental particles (SM), neutrinos are massless particles. Such a conclusion is challenged by the number density measurements of the solar and atmospheric neutrinos, pointing to an oscillatory behavior of the three flavors. Oscillation experiments are insensitive to the absolute mass scale of the neutrinos, but useful to probe their squared mass differences. The latter results from atmospheric neutrinos point to $|\Delta m_{32}^{2}|\approx 2.4\times 10^{-3}\;\text{eV}^{2}$, while from solar neutrinos one obtains $|\Delta m_{21}^{2}|\approx 7.4\times 10^{-5}\;\text{eV}^{2}$ [31]. This configuration enables at least two non-relativistic species of neutrinos. Lower limits on the sum of their masses can also be estimated: $\sum m_{\nu}>0.06\;\text{eV}$ for the Normal Hierarchy schemes (NH), while the Inverted Hierarchy (IH) requires $\sum m_{\nu}>0.10\;\text{eV}$.
Back to the cosmological perspective, neutrinos have a well-marked role in the expansion of the cosmos. They act as radiation in the very early universe, later changing to a hot dark matter behavior after the non-relativistic transition. At the perturbative level, the neutrinos dispersion acts to suppress the growth of perturbation at small scales, characterized by their free-streaming and, ultimately, by the sum of their masses. One of the current tightest constraints, $\sum m_{\nu}<0.07\;\text{eV}$ at 95% confidence level (C.L), comes from the combination of CMB and DESI BAO data within the $\Lambda$CDM framework, showing a clear preference for the NO [29, 32] (see also [33]). More recently, the upper limit $\sum m_{\nu}<0.042$ eV was obtained using CMB combined with DESI BAO, galaxy cluster measurements, and a $H_{0}$ prior from SH0ES [34]. These bounds are substantially relaxed under a DDE framework modeled by CPL, yielding $\sum m_{\nu}<0.11\;\text{eV}$ at 95% C.L [29](see also [35]). Therefore, the current cosmological limits for the $\Lambda$CDM are on the verge of excluding both inverted and normal ordering, creating tension with the results of particle physics experiments [36, 34].
@@ III Analysis and Results
To obtain constraints on the cosmological parameters for the DDE models above, we perform a Monte Carlo Markov Chain (MCMC) numerical analysis using Planck 2018 temperature and polarization data, including low- and high-multipole measurements of temperature and polarization, combined with lensing (plikHM+TTTEEE+lowl+lowE+lensing) [38], along with Pantheon+ catalog SN data [39, 40] and geometric BAO information from the DESI collaboration data release two [29]. For this, we implement the parameterizations in the Boltzmann solver code Cosmic Linear Anisotropy Solving System (CLASS) [41], in which we interfaced with the Cobaya code to perform the MCMC statistical analysis [42, 43]. For the purpose of constraining the neutrino mass, we assume the standard degenerate mass spectrum, in which the three massive neutrino species each carry one third of the total mass, i.e., $\sum m_{\nu}=3(m_{\nu,1}=m_{\nu,2}=m_{\nu,3})$. See [44] for a detailed discussion on the choice of neutrino mass hierarchy in cosmological analyses. In our case, we also imposed a prior that the total neutrino mass must satisfy $\sum m_{\nu}>0$.
Figure 1: Constraints on $\sum m_{\nu}$ from Planck 2018 temperature, polarization and lensing together with Pantheon+ SN and DESI DR2 data assuming JBP, BA, and Thawing parameterizations. The dashed vertical line represents the inverted hierarchy $\sum m_{\nu}>0.10$ eV. For comparison, we also show the CPL prediction (black line).
Figure 1 shows the marginalized posterior distribution for $\sum m_{\nu}$ across the models explored in this work. For comparison, see Fig. 7 of [45], which presents results from the DESI collaboration using DESI DR2 BAO + CMB + Pantheon+ data. In their analysis of the CPL model, the posterior peaks at the prior boundary $\sum m_{\nu}=0$, suggesting that the data favor non-physical values of $\sum m_{\nu}<0$. This is not the case in our analysis: all models show a peak at $\sum m_{\nu}>0$, with only a small portion of the posterior excluded due to prior choice $\sum m_{\nu}>0$. It is important to note that DESI uses Planck combined with Data Release 6 of the Atacama Cosmology Telescope (ACT), while in our analysis we use CMB data from the Planck baseline likelihood only.
@@
CPL $0.048^{+0.076}_{-0.049}$ eV $0.0$
BA $0.050^{+0.079}_{-0.051}$ eV $-0.64$
JBP $0.036^{+0.061}_{-0.037}$ eV $-0.60$
TWA $0.027^{+0.045}_{-0.027}$ eV $2.66$
Table 1: $2\sigma$ constraints on the sum of neutrino masses $\sum m_{\nu}$ and the Bayes factor, $\Delta\ln{B}$, for all models discussed in the text. Negative values of $\Delta\ln{B}$ favors the reference model (CPL).
In Figure 2, we show the $w_{0}$$w_{a}$ planes for JBP (upper panel) and BA parameterizations (bottom panel). When treating neutrino mass as a free parameter, the JBP model is consistent with the $\Lambda$CDM scenario within $2\sigma$, while the BA model lies nearly outside the $2\sigma$ boundary. Following the procedure of [20], we also project the $(\alpha,\beta)$ values from the thawing model into the $(w_{0},w_{a})$ parameter space by matching the expansion history $H(z)$. The shaded regions in both panels of Fig. 2 represent the $\alpha$$\beta$ values within the full prior range considered in the analysis: $\alpha\in[0,1]$ and $\beta\in[0,15]$. This shows that the phantom region in the $w_{0}$$w_{a}$ plane for both models is fully compatible with the scenario of thawing quintessence. The star markers correspond to the best-fit values, $\alpha=0.15$ and $\beta=7.84$, projected into the $w_{0}$$w_{a}$ parameter spaces for both models. These best-fit projections lie within the $1\sigma$ and $2\sigma$ confidence regions of the JBP and BA, respectively.
@@ IV Final Remarks
Our findings in this work highlight several key aspects. First, unlike previous results reported by the DESI collaboration [48], where the posterior for $\sum m_{\nu}$ in the CPL model peaks at the prior edge of $\sum m_{\nu}=0$, which indicates a preference for non-physical negative neutrino mass values, all models in our analysis peak at positive values. This suggests that Planck data when combined with ACT (used in DESI’s analysis) may significantly shift the results toward this negative preference. In addition, the thawing parameterization provides the strongest constraint on the neutrino mass, $\sum m_{\nu}<0.071$ eV at 95% C.L., comparable to that of $\Lambda$CDM and sufficient to exclude the inverted hierarchy at more than $3\sigma$. The JBP model also shows a mild preference for the normal hierarchy over the inverted, with an upper bound of $\sum m_{\nu}<0.096$ eV at 95% C.L. We find that the JBP model is consistent with $\Lambda$CDM within $2\sigma$ while the BA model sits near the outer limit of $2\sigma$. Both models exhibit a shift toward $\Lambda$CDM when the neutrino mass is included in the analysis as a free parameter. This is significant, as it weakens the apparent preference for DDE models seen in analyses where the total neutrino mass is fixed at $\sum m_{\nu}=0.06$ eV.
@@
Finally, we studied the redshift evolution of the dark energy equation of state. Although they all remain in the quintessence regime at low redshifts, their behaviors diverge at higher redshifts. The thawing model stands out because of its smooth transition from a cosmological constant-like state ($w=-1$) at high redshift to a quintessence-like behavior near the present epoch. Overall, our results provide evidence that thawing quintessence models are not only fully compatible with widely used phenomenological parameterizations when projected in the same parameter space, but also offer tighter bounds on the neutrino mass. The thawing model also shows a strong preference for the normal mass ordering and for positive, nonzero values of $\sum m_{\nu}$.

### v3

@@ Abstract
Recent analyses have shown that a dynamic dark energy modeled by the CPL parameterization of the dark energy equation of state (EoS) can ease constraints on the total neutrino mass compared to the standard $\Lambda$CDM model. This helps reconcile cosmological and particle physics measurements of $\sum m_{\nu}$. In this study, we investigate the robustness of this effect by assessing the extent to which the CPL assumption influences the results. We examine how alternative EoS parameterizations - such as Barboza-Alcaniz (BA), Jassal-Bagla-Padmanabhan (JBP), and a physically motivated thawing parameterization that reproduces the behavior of various scalar field models - affect estimates of $\sum m_{\nu}$. Although both the BA and JBP parameterizations relax the constraints similarly to the CPL model, the JBP parameterization still excludes the inverted neutrino mass hierarchy at $\sim 2.1\;\sigma$ with $\sum m_{\nu}<0.096$ eV. The thawing parameterization excludes the inverted hierarchy at $\sim 3.3\sigma$ and yields tighter constraints, comparable to those of the $\Lambda$CDM model, with $\sum m_{\nu}<0.071$ eV. Finally, we show that the thawing model can be mapped into the BA and JBP $w_{0}$$w_{a}$ parameter space, with the apparent preference for the phantom regime actually supporting quintessence (non-phantom) models.
@@ I Introduction
In the context of General Relativity, two major frameworks have emerged as alternatives to model the Universe’s accelerated expansion (see e.g. [5, 6, 7, 8, 9, 10, 11]): the scalar field approach, exemplified by quintessence models, and phenomenological parameterizations of the dark energy equation of state (EoS), such as the widely used Chevallier–Polarski–Linder (CPL), Barboza–Alcaniz (BA), and Jassal–Bagla–Padmanabhan (JBP) forms [12, 13, 14, 15, 16]. In both frameworks, the dark energy EoS is a dynamical quantity, $w(z)$.
@@
Beyond the DDE discussion, the late-time universe is plagued by further uncertainties related to the background and perturbations evolution, and the neutrino physics is of particular interest in this picture. In the standard scenario, outlined by the Standard Model of fundamental particles (SM), neutrinos are massless particles. Such a conclusion is challenged by the number density measurements of the solar and atmospheric neutrinos, pointing to an oscillatory behavior of the three flavors. Oscillation experiments are insensitive to the absolute mass scale of the neutrinos, but useful to probe their squared mass differences. The latter results from atmospheric neutrinos point to $|\Delta m_{32}^{2}|\approx 2.4\times 10^{-3}\;\text{eV}^{2}$, while from solar neutrinos one obtains $|\Delta m_{21}^{2}|\approx 7.4\times 10^{-5}\;\text{eV}^{2}$ [31]. This configuration enables at least two non-relativistic species of neutrinos. Lower limits on the sum of their masses can also be estimated: $\sum m_{\nu}>0.06\;\text{eV}$ for the Normal Ordering schemes (NO), while the Inverted Ordering (IO) requires $\sum m_{\nu}>0.10\;\text{eV}$.
Back to the cosmological perspective, neutrinos have a well-marked role in the expansion of the cosmos. They act as radiation in the very early universe, later changing to a hot dark matter behavior after the non-relativistic transition. At the perturbative level, the neutrinos’ dispersion acts to suppress the growth of perturbations at small scales, characterized by their free-streaming and, ultimately, by the sum of their masses. Increasing $\sum m_{\nu}$ also changes the amplitude of the distance measurements in late-time Universe. In particular, the discrepancy between the values of $H_{0}r_{d}$ measured by CMB observations and BAO measurements can be mitigated by reducing $\sum m_{\nu}$ in the context of $\Lambda$CDM [32]. One of the current tightest constraints, $\sum m_{\nu}<0.07\;\text{eV}$ at 95% confidence level (C.L), comes from the combination of CMB and DESI BAO data within the $\Lambda$CDM framework, showing a clear preference for the NO [29, 33] (see also [34]). More recently, the upper limit $\sum m_{\nu}<0.042$ eV was obtained using CMB combined with DESI BAO, galaxy cluster measurements, and a $H_{0}$ prior from SH0ES [35]. Therefore, the current cosmological limits for the $\Lambda$CDM are on the verge of excluding both inverted and normal ordering, creating tension with the results of particle physics experiments [36, 35].
These bounds are substantially relaxed under a DDE framework modeled by CPL, yielding $\sum m_{\nu}<0.11\;\text{eV}$ at 95% C.L [29](see also [37]). When expanding the cosmological model to a dynamical dark energy scenario, larger values of $\sum m_{\nu}$ are allowed by decreasing the value of $w(z)$, since these two parameters are anti-correlated [38, 39]. The degeneracy of $\sum m_{\nu}$ and $w(z)$ is partially broken when CMB, BAO and supernova data are considered concomitantly.
@@ III Analysis and Results
To obtain constraints on the cosmological parameters for the DDE models above, we perform a Monte Carlo Markov Chain (MCMC) numerical analysis using Planck 2018 temperature and polarization data, including low- and high-multipole measurements of temperature and polarization, combined with lensing (plikHM+TTTEEE+lowl+lowE+lensing) [41] [^1] It is important to note that DESI uses Planck combined with Data Release 6 of the Atacama Cosmology Telescope (ACT), while in our analysis we use CMB data from the Planck baseline likelihood only., along with Pantheon+ catalog SN data [42, 43] and geometric BAO information from the DESI collaboration data release two [29]. For this, we implement the parameterizations in the Boltzmann solver code Cosmic Linear Anisotropy Solving System (CLASS) [44], in which we interfaced with the Cobaya code to perform the MCMC statistical analysis [45, 46]. For the purpose of constraining the neutrino mass, we assume the standard degenerate mass spectrum, in which the three massive neutrino species each carry one third of the total mass, i.e., $\sum m_{\nu}=3(m_{\nu,1}=m_{\nu,2}=m_{\nu,3})$. See [47] for a detailed discussion on the choice of neutrino mass hierarchy in cosmological analyses. In our analysis, we also imposed a prior that the total neutrino mass must satisfy $\sum m_{\nu}>0$, with $\sum m_{\nu}\in[0,5]$. For the EoS parameters our priors are, $w_{0}\in[-3,1]$, $w_{a}\in[-3,2]$, $\alpha\in[0,1]$ and $\beta\in[0,15]$.
Figure 1: Upper bounds on $\sum m_{\nu}$ from Planck 2018 temperature, polarization and lensing together with Pantheon+ SN and DESI DR2 data assuming JBP, BA, and Thawing parameterizations. The dashed vertical line represents the inverted hierarchy $\sum m_{\nu}>0.10$ eV. For comparison, we also show the CPL prediction (black line).
Figure 1 shows the marginalized posterior distribution for $\sum m_{\nu}$ across the models explored in this work. For comparison, see Fig. 7 of [48], which presents results from the DESI collaboration using DESI DR2 BAO + CMB + Pantheon+ data. In their analysis of the CPL model, the posterior peaks at the prior boundary $\sum m_{\nu}=0$, suggesting that the data favor non-physical values of $\sum m_{\nu}<0$. This is also true in our analysis: all models show a peak at $\sum m_{\nu}=0$. This suggests that regardless of the DDE model, when including DESI data, there is still a preference outward $\sum m_{\nu}\leq 0$ values.
@@
CPL $<0.123$ eV $0.0$
BA $<0.127$ eV $-0.64$
JBP $<0.096$ eV $-0.60$
TWA $<0.0709$ eV $2.66$
Table 1: $2\sigma$ upper bounds on the sum of neutrino masses $\sum m_{\nu}$ and the Bayes factor, $\Delta\ln{B}$, for all models discussed in the text. Negative values of $\Delta\ln{B}$ favors the reference model (CPL).
In Figure 2, we show the $w_{0}$$w_{a}$ planes for JBP (upper panel) and BA parameterizations (bottom panel). When treating neutrino mass as a free parameter, the JBP model is consistent with the $\Lambda$CDM scenario within $2\sigma$, while the BA model lies nearly outside the $2\sigma$ boundary. Following the procedure of [20], we also project the $(\alpha,\beta)$ values from the thawing model into the $(w_{0},w_{a})$ parameter space by matching the expansion history $H(z)$. The shaded regions in both panels of Fig. 2 represent the $\alpha$$\beta$ values within the full prior range considered in the analysis. This shows that the phantom region in the $w_{0}$$w_{a}$ plane for both models is fully compatible with the scenario of thawing quintessence. The star markers correspond to the best-fit values, $\alpha=0.15$ and $\beta=7.84$, projected into the $w_{0}$$w_{a}$ parameter spaces for both models. These best-fit projections lie within the $1\sigma$ and $2\sigma$ confidence regions of the JBP and BA, respectively.
@@ IV Final Remarks
Our findings in this work highlight several key aspects. First, regardless of the underlying cosmological model, be it $\Lambda$CDM or a DDE scenario, the posterior distribution for $\sum m_{\nu}$, when DESI data are included, consistently peaks at the lower prior boundary, $\sum m_{\nu}=0$. This behavior indicates that the data would prefer values of $\sum m_{\nu}$ that are negative, an unphysical regime excluded by the prior. Consequently, the derived upper bounds on $\sum m_{\nu}$ in this context are largely driven by the prior choice. The thawing parameterization provides the strongest constraint on the neutrino mass, $\sum m_{\nu}<0.071$ eV at 95% C.L., comparable to that of $\Lambda$CDM and sufficient to exclude the inverted hierarchy at more than $3\sigma$. The JBP model also shows a mild preference for the normal hierarchy over the inverted, with an upper bound of $\sum m_{\nu}<0.096$ eV at 95% C.L. We find that the JBP model is consistent with $\Lambda$CDM within $2\sigma$ while the BA model sits near the outer limit of $2\sigma$. Both models exhibit a shift toward $\Lambda$CDM when the neutrino mass is included in the analysis as a free parameter. This is significant, as it weakens the apparent preference for DDE models seen in analyses where the total neutrino mass is fixed at $\sum m_{\nu}=0.06$ eV.
@@
Finally, we studied the redshift evolution of the dark energy equation of state. Although they all remain in the quintessence regime at low redshifts, their behaviors diverge at higher redshifts. The thawing model stands out because of its smooth transition from a cosmological constant-like state ($w=-1$) at high redshift to a quintessence-like behavior near the present epoch. Overall, our results provide evidence that thawing quintessence models are not only fully compatible with widely used phenomenological parameterizations when projected in the same parameter space, but also offer tighter bounds on the neutrino mass. The thawing model also shows a strong preference for the normal mass ordering.